OpenStudy (anonymous):

sin(x+10deg)=cos3x then x=?

OpenStudy (anonymous):

\[\sin(x+10)=\cos(3x)\] \[\sin(a+b)=\sin(a)\cos(b)+\sin(b)\cos(a)\]

OpenStudy (anonymous):

\[\tan(x)=\frac{\sin(x)}{\cos(x)}\]This is probably all you'll need

OpenStudy (anonymous):

but why it became tanx= sinx/cosx ?

OpenStudy (anonymous):

No, that's just the equation for tan

OpenStudy (anonymous):

how will i distribute it? :)